OpenStudy (anonymous):

Which point is a relative minimum for the function y = -1 + csc(x)? a) (3pi/2, 0) b) (3pi/2, -1) c) (pi/2, 0) d) (pi/2, -1)

OpenStudy (anonymous):

y'=??

OpenStudy (anonymous):

idk im super confused

OpenStudy (anonymous):

so take the derivative first

OpenStudy (anonymous):

D[csc(x)]=-csc(x)cot(x)

OpenStudy (anonymous):

next lets simplify it a little bit. \[\frac{ -1 }{ \sin(x) }*\frac{\cos(x)}{\sin(x)}\]

OpenStudy (anonymous):

\[\frac{ -\cos(x) }{ \sin^2 (x) }\]

OpenStudy (anonymous):

set this equal to 0

OpenStudy (anonymous):

are you with me?

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

okay just tell me where you got lost

OpenStudy (anonymous):

D[csc(x)]=-csc(x)cot(x) im not understanding what that is

OpenStudy (anonymous):

D stands for derivative

OpenStudy (anonymous):

That is the derivative of csc(x)

OpenStudy (anonymous):

^^ bebong got it